The mathematics club of RAD Senior High School organized a raffle where the prize is the latest edition of a certain brand of the calculator. the raffle aims to raise funds for their club. they were able to sell 1000 tickets that cost 5 pesos each. if they spent 3500 on the prize, what will be the expected gain of a student who will buy one ticket?

1
Expert's answer
2021-05-02T08:51:25-0400

E(X)="\sum{ }"X"\times"P(X)

There's a probability of 1/1000 to win 3500 pesos and 999/1000 to loose 5 pesos

E(X) = "\sum" (3500"\times\frac{1}{1000}+(-5)\times\frac{999}{1000}") = -1.495